### v3

# Exploring Test-Time Adaptation for Object Detection in Continually Changing Environments

Abstract.
Real-world application models are commonly deployed in dynamic environments, where the target domain distribution undergoes temporal changes. Continual Test-Time Adaptation (CTTA) has recently emerged as a promising technique to gradually adapt a source-trained model to continually changing target domains. Despite recent advancements in addressing CTTA, two critical issues remain: 1) Fixed thresholds for pseudo-labeling in existing methodologies generate low-quality pseudo-labels, as model confidence varies across categories and domains; 2) Stochastic parameter restoration methods for mitigating catastrophic forgetting fail to effectively preserve critical information due to their intrinsic randomness. To tackle these challenges for detection models in CTTA scenarios, we present CTAOD, featuring three core components. Firstly, the object-level contrastive learning module extracts object-level features for contrastive learning to refine the feature representation in the target domain. Secondly, the adaptive monitoring module dynamically skips unnecessary adaptation and updates the category-specific threshold based on predicted confidence scores to enable efficiency and improve the quality of pseudo-labels. Lastly, the data-driven stochastic restoration mechanism selectively reset inactive parameters with higher possibilities, ensuring the retention of essential knowledge. We demonstrate the effectiveness of CTAOD on four CTTA object detection tasks, where CTAOD outperforms existing methods, especially achieving a 3.2 mAP improvement and a 20% increase in efficiency on the Cityscapes-to-Cityscapes-C CTTA task. The code will be released.
copyright: none
1. Introduction
Figure 1. Motivation. The source-trained model adapts to dynamic environments without source data. (1) AM: The bar chart shows mean model scores varying across categories and domains. Unlike existing methods that continually adapt to all test data using fixed thresholds, CTAOD dynamically skips unnecessary adaptations and updates the category-specific threshold. (2) DSR: The bottom illustrates the neuron reset process, where darker purple indicates higher activity, and blue represents source knowledge. While existing methods restore parameters randomly, CTAOD resets inactive parameters with higher possibilities.
Deep learning models have demonstrated immense potential across various data modalities such as image (He et al., 2016; Dosovitskiy et al., 2020), audio (Aytar et al., 2016), and language (Devlin et al., 2018). However, these models experience pronounced degradation in performance when confronted with training data (i.e., source domain) and testing data (i.e., target domain) originating from disparate distributions. This phenomenon, commonly referred to as distribution shifts, poses a significant challenge (Geirhos et al., 2018; Hendrycks and Dietterich, 2019).
In such a scenario, unsupervised domain adaptation (UDA) becomes crucial, which typically involves aligning the source and target data distributions, thereby mitigating the impact of distribution shifts (Ganin et al., 2016; Tzeng et al., 2017). Still, UDA falls short by necessitating access to source data, which is often inaccessible due to privacy constraints or data transmission barriers (Huang et al., 2021).
This limitation catalyzes the exploration of Test-Time Adaptation (TTA) where the source-trained model directly adapts toward unlabeled test samples encountered during evaluation in an online manner, without the reliance on the source data (Wang et al., 2020). Nonetheless, the TTA methods, which assume a static target domain, face a more challenging and realistic problem, as real-world systems work in non-stationary environments. For example, a vehicle may encounter various continuous environmental changes such as fog, night, rain, and snow during its journey. Existing TTA methods are vulnerable to catastrophic forgetting of source knowledge and error accumulation when adaptation faces more than one distribution shift (Wang et al., 2022).
Recently, Wang et al. (2022) introduce CoTTA by applying stochastic parameters restoration to mitigate catastrophic forgetting in Continual Test-Time Adaptation (CTTA) scenarios, where the model is continually adapted to sequences of target domains. Although randomly resetting parameters partially helps alleviate the forgetting of source knowledge (Wang et al., 2022; Zhu et al., 2023), its randomness may also contribute to losing crucial knowledge specific to the current domain (refer to Figure 1). Additionally, existing CTTA methods normally utilize pseudo-labeling through a fixed threshold for self-supervision (Wang et al., 2022; Döbler et al., 2023). Given that model confidence may vary across categories and domains (refer to Figure 1), employing a uniformly fixed threshold could exclude high-quality pseudo-labels while incorporating incorrect ones. Furthermore, these methods adapt to every new incoming data, resulting in computational inefficiency and potential performance degradation. Low-quality pseudo-labels and unstable adaptation to noisy data lead to error accumulation, as they provide negative feedback to the model.
Therefore, despite recent efforts to handle CTTA, two significant challenges persist: Current self-training-based methods suffer from low-quality pseudo-labels, leading to error accumulation (Challenge 1); Continual adaptation to dynamic environments struggles to effectively retain valuable knowledge about the current domain while mitigating the forgetting of source knowledge (Challenge 2).
To tackle these challenges, we introduce CTAOD (Continual Test-time Adaption for Object Detection). Aligning with previous CTTA works for robust adaptation (Wang et al., 2022), CTAOD is constructed upon the mean-teacher framework (Tarvainen and Valpola, 2017) where the teacher model is an exponential moving average of the student model. Particularly, CTAOD comprises Object-level Contrastive Learning (OCL), Adaptive Monitoring (AM), and Data-driven Stochastic Restoration (DSR).
More specifically, OCL extracts features based on proposals generated by the Region Proposal Network (RPN), which provides multiple cropped views around the object at different locations and scales. Subsequently, Contrastive Learning (CL) loss is applied on RPN cropped views to guide the model to encourage similar object instances to remain close while pushing dissimilar ones apart. The OCL is well integrated into the mean-teacher paradigm as a drop-in enhancement to acquire more fine-grained and localized feature representation. Furthermore, to make adaptation more efficient and stable and to improve the quality of pseudo-labels, i.e., addressing Challenge 1, we design an AM module to decide whether to pause or resume the adaptation and dynamically adjust the category-specific thresholds, based on the predicted confidence scores. The dynamic nature of the AM method makes it better suited to address the effects of continuously changing distributions. Finally, the DSR mechanism is proposed to reset inactive parameters with a higher possibility than active ones by employing the gradients of parameters as weights for a random mask matrix. DSR not only helps prevent forgetting but preserves important information, i.e., addressing Challenge 2. On the other hand, its randomness also allows potentially falsely activated parameters to be reset, thereby making the model more stable.
We demonstrate the effectiveness of our proposed approach on four continual test-time adaptation benchmark tasks for object detection. These datasets are associated with synthetic and real-world distribution shifts in the short-term and long-term adaption, i.e. Cityscapes (Cordts et al., 2016), Cityscapes-C (Hendrycks and Dietterich, 2019), SHIFT (Sun et al., 2022), and ACDC (Sakaridis et al., 2021). Experimental results indicate that our method significantly improves performance over existing methods, with gains of up to 3.2 mAP and 20% in computational efficiency.
Our main contributions are summarized as follows:
  • This study introduces CTAOD, which pioneers exploring CTTA for detection models. Specifically, we propose to leverage object-level features for contrastive learning to refine feature representation in CTTA object detection.
  • To address the two challenges in CTTA, our proposed CTAOD enables dynamic skipping and category-specific threshold updates based on predicted scores and reset the inactive parameter with higher possibilities.
  • Empirical experiments demonstrate that CTAOD surpasses existing methods and facilitates efficient short-term and long-term adaptation in dynamic environments.
2. RELATED WORK
2.1. Source-Free Domain Adaptation
UDA tackles the inter-domain divergence by aligning the distributions of source and target data (Ganin et al., 2016; Luo et al., 2020). Despite its effectiveness, the limitation of UDA lies in its requirement for access to the source domain data, which often raises concerns regarding data privacy and transmission efficiency. As a result, Source-Free Domain Adaptation (SFDA) received extensive research attention in object detection, where the source-trained detector is re-trained on the target training data (VS et al., 2023b, a; Li et al., 2021) before evaluation. For instance, MemCLR (VS et al., 2023a) employs a cross-attention-based memory bank for CL, while IRG (VS et al., 2023a) utilizes the object relations with instance relation graph network to explore the SFDA setting for object detection without the source data. However, the standard SFDA setting requires prior knowledge of the target domain, which is impractical in most real-world applications.
2.2. Test-Time Adaptation
TTA adapts the source-trained model to the target test data during inference time without access to the source data. Since both TTA and SFDA involve adapting the source-trained model to the unlabeled target data without utilizing source data, some works also refer to TTA as SFDA (Wang et al., 2022; Brahma and Rai, 2023). In this paper, we distinguish between TTA and SFDA based on evaluation protocol, although they can be transformed into each other in experiments. Furthermore, TTA methods improve the model performance under distribution shift commonly through pseudo-labeling (Sinha et al., 2023; Iwasawa and Matsuo, 2021), batchnorm statistics updating (Hu et al., 2021; You et al., 2021), or entropy regularization (Wang et al., 2020; Iwasawa and Matsuo, 2021) during testing. For example, Tent (Wang et al., 2020) updates the batchnorm parameters with entropy minimization and demands a large batch size for optimization during test-time adaptation, which is unsuitable for real-time detection model deployment where images are processed sequentially. TTA methods assume a static target domain where the target data come from a single domain. However, the target distribution may exhibit a continual shift over time in practical scenarios.
2.3. Continual Test-time Adaptation
Recently, Wang et al. (2022) introduce CoTTA, marking the first work tailored to the demands of CTTA by adapting a pre-trained model to sequences of domains without source data. Subsequently, research efforts have been dedicated to exploring CTTA, primarily focused on classification (Döbler et al., 2023; Brahma and Rai, 2023; gan2023decorate) and segmentation (Song et al., 2023; Ni et al., 2023; Liu et al., 2023) tasks. For instance, similar to CoTTA, Zhu et al. (2023) apply a stochastic reset mechanism to prevent forgetting in the CTTA medical segmentation task. In contrast, PETAL (Brahma and Rai, 2023) utilizes the Fisher Information Matrix (FIM) as a metric of parameter importance to reset only the most irrelevant parameters. Nevertheless, these two methods are sub-optimal, as the randomness might lead to losing essential information, and pure data-driven restoration may retain false active parameters resulting from noise.
Moreover, the mean-teacher framework (Tarvainen and Valpola, 2017) serves as a base architecture for most CTTA works (Wang et al., 2022; Gong et al., 2022), where the teacher generates pseudo-labels via a fixed threshold to supervise the student. Nonetheless, these methods suffer from low-quality pseudo-labels with a uniform threshold since the model confidence varies across categories and domains. Furthermore, Wang et al. (2024) design a dynamic thresholding technique to update the threshold in a batch for classification tasks, requiring a large batch size. However, this approach is unsuitable for object detection where smaller batch sizes are preferred for computational efficiency. To improve efficiency, yoo2024and presents a CTTA approach for object detection to update specific parts of the model in specific adaptation, which utilizes source statistics. Although it effectively improved efficiency, its performance is limited. Additionally, Gan et al. (2023) presents a cloud-device collaborative continual adaptation paradigm in object detection by aligning source and target distribution. Nevertheless, the reliance on source data in both methods restricts their practical applicability. Therefore, a gap still exists in exploring the CTTA in object detection to improve the quality of pseudo-label (Challenge 1) and effectively reset noisy neurons (Challenge 2), without relying on any source data.
3. METHOD
3.1. Preliminary
3.1.1. Problem Statement
Given a sequence of domain $D=\{D^{0},D^{1},...,D^{n}\}$, we define $D^{0}$ as the source domain and the subsequent domains as the target domain. The objective of CTTA is to enhance the performance of the model $f_{\theta_{0}}(x)$, where the parameters $\theta_{0}$ are pre-trained on source data $x^{D^{0}},y^{D^{0}}$ from $D^{0}$, in a continually changing target domain during inference time without using source data. For simplicity, we denote the data without the superscript about the specific domain in the following discussion. At time step $t$, unlabeled target data $x_{t}$ is provided sequentially, following the domain group order. The model is required to make a prediction $\hat{y}_{t}=f_{\theta_{t-1}}(x_{t})$ using the parameters $\theta_{t-1}$ which have been updated based on previous target data $x_{1},...,x_{t-1}$. Subsequently, $\hat{y}_{t}$ serves as the evaluation output at time step t, and the model will adapt itself toward $x_{t}$ as $\theta_{t-1}\rightarrow\theta_{t}$, which will only influence future inputs $x_{t+n}$.
Figure 2. Overview. CTAOD follows the mean-teacher framework, featuring: (1) The OCL module compares region of interest features extracted from feature maps of both networks based on teacher proposals for contrastive learning; (2) The AM module dynamically skips unnecessary adaptation and adjusts category-specific thresholds based on prediction scores; (3) The DSR reset the inactive parameter with higher possibilities based on the FIM.
3.1.2. Mean-Teacher Framework
In alignment with previous CTTA works (Wang et al., 2022; Gong et al., 2022), we build our method based on the mean-teacher framework (Tarvainen and Valpola, 2017). Specifically, both networks are first initialized with the source-trained model. The teacher model produces the pseudo-labels for the unlabeled target data to supervise the student model. While the parameters of the student model are optimized via gradient descent, the teacher model is updated following an Exponential Moving Average (EMA) strategy based on the student model. Formally, this process can be expressed as follows:
(1) $\mathcal{L}_{pl}(x_{t})=\mathcal{L}_{rpn}(x_{t},\hat{y}_{t})+\mathcal{L}_{rcnn}(x_{t},\hat{y}_{t}),$
(2) $\theta_{t}^{std}\leftarrow\theta_{t-1}^{std}+\gamma\frac{\partial(\mathcal{L}_{pl}(x_{t}))}{\partial\theta_{t-1}^{std}},$
(3) $\theta_{t}^{tch}\leftarrow\alpha\theta_{t-1}^{tch}+(1-\alpha)\theta_{t}^{std},$
where $x_{t}$ and $\hat{y}_{t}$ denote the unlabeled target data and corresponding pseudo-labels. $\theta^{std}_{t}$ and $\theta^{tch}_{t}$ symbolize the parameters of student and teacher networks at time step $t$. Moreover, the supervision loss $\mathcal{L}_{pl}$ in Faster-RCNN (Ren et al., 2015), which consists of the RPN loss $\mathcal{L}_{rpn}$ and RCNN loss $\mathcal{L}_{rcnn}$, is utilized for pseudo-labeling. Additionally, $\gamma$ represents the student’s learning rate, and the EMA rate is denoted by $\alpha$. Hence, the teacher can be regarded as an ensemble of historical students, providing stable supervision. However, this framework encounters error accumulation and catastrophic forgetting in dynamic environments. Therefore, we design CTAOD to fit in the CTTA setting.
3.2. Proposed Method
3.2.1. Overview
Figure 2 presents a concise overview of CTAOD, consisting of Object-level Contrastive Learning (OCL), Adaptive Monitoring (AM), and Data-driven Stochastic Restoration (DSR). Inspired by UDA object detection methods (Li et al., 2022; Liu et al., 2021), we adopt the Weak-Strong augmentation to enable the teacher model to generate reliable pseudo-labels without being affected by heavy augmentation. Specifically, the teacher model receives input with weak augmentations while we input student networks with strong enhanced images. The overall loss of the student model for gradient descent is defined as:
(4) $\mathcal{L}_{all}=\mathcal{L}_{pl}(x_{t})+\lambda\mathcal{L}_{cl}(x_{t})+\mu\mathcal{L}_{kl}(x_{t}),$
where $\mathcal{L}_{cl}(x_{t})$ denotes the contrastive learning (CL) loss, to be elaborated later. The $\mathcal{L}_{kl}$ represents the Kullback-Leibler (KL) Divergence (Kullback and Leibler, 1951) loss, used to quantify the distinction between two probability distributions. The parameters $\lambda$ and $\mu$ are corresponding weights of the losses. The KL loss is defined as:
(5) $\mathcal{L}_{kl}\left(P\parallel Q\right)=\sum_{x\in\mathcal{X}}P(x)\mathrm{~{}}\log\left(\frac{P(x)}{Q(x)}\right),$
where $P(x)$ and $Q(x)$ represent different distribution. We employ the KL divergence loss to encourage the student model to approximate the teacher model closely.
3.2.2. Object-Level Contrastive Learning
SimCLR is a widely used CL approach for self-supervised visual representation learning (Chen et al., 2020). SimCLR learns high-quality feature representations from differently augmented views of the same image. Designed for classification tasks, it requires large batch sizes to ensure sufficient positive and negative pairs. Consequently, SimCLR is not well-suited for object detection, where images contain multiple instances and large batch sizes demand significant computational resources.
Motivated by SimCLR (Chen et al., 2020), we present an OCL module to extract teacher and student features for CL based on proposals generated from the RPN. This strategy provides multiple cropped views around the object instance, eliminating the need for large batch sizes and ensuring computational efficiency for online detection model updates. Specifically, given a weakly augmented image $f_{weak}(x_{t})$ at time step t, the teacher produces ROI proposals $p_{t}=\{p_{1},...,p_{l}\}$ via region proposal network. We then apply RoIAlign (He et al., 2017), a pooling operation for ROI, to extract corresponding teacher and student object-level features $F_{t}^{T}=\{f_{i}^{T}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ and $F_{t}^{S}=\{f_{i}^{S}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ based on the feature map from the backbone, respectively. The features associated with the same proposal are considered positive pairs, otherwise negative pairs. we employ the CL loss to these features $F_{t}^{T}$ and $F_{t}^{S}$ derived from weak and strong enhanced images respectively by minimizing:
(6) $\mathcal{L}_{cl}(x_{t})=\frac{1}{l}\sum_{i=1}^{l}-\log\frac{\exp(f_{i}^{T}\cdot f_{i}^{S}/\tau)}{\sum_{j=1}^{l}\exp(f_{i}^{T}\cdot f_{j}^{S}/\tau)},$
where $\tau>0$ is the temperature, $l$ indicates the number of the features, and $f_{i}^{T}$ and $f_{i}^{S}$ denote the features of two different augmentations of the same object, serving as the positive pair. This strategy encourages the model to learn fine-grained and localized representations on the target domain, without relying on accurate pseudo-labels.
3.2.3. Adaptive Monitoring
The AM module monitors the model’s status using the predicted scores, featuring dynamic skipping and dynamic threshold. The former halts unnecessary adaptations to save computational resources, while the latter updates the category-specific threshold.
As shown in Figure 2, the teacher network first make predictions $\hat{y}_{t}=f_{\theta_{t-1}}(f_{wek}(x_{t}))$ for the weakly enhanced images. Subsequently, We compute the mean scores of all category $\overline{l}_{t}$ in teacher’s predictions and maintain its exponentially moving average $\overline{l}_{t}^{ema}$ through:
(7) $\overline{l}_{t}^{ema}\leftarrow\beta_{s}\cdot\overline{l}_{t-1}^{ema}+(1-\beta_{s})\cdot\overline{l}_{t},$
where $\beta_{s}$ represents the update rate. We introduce a new index $\frac{\overline{l}_{t}}{\overline{l}_{t}^{ema}}$. Adaptation resumes when the $\frac{\overline{l}_{t}}{\overline{l}_{t}^{ema}}$ lies within the range $(\frac{1}{\delta_{s}},\delta_{s})$, indicating stability, and pauses when it deviates beyond this range, suggesting changes in data characteristics. This dynamic skipping strategy effectively prevents unstable and noisy adaptation, thus conserving computational resources and mitigating error accumulation.
For dynamic threshold, the thresholds for each category are initialized with the same value $\delta_{0}$ at first, which is updated every iteration by:
(8) $\delta_{t}^{c}\leftarrow\beta_{t}\cdot\delta_{t-1}^{c}+(1-\beta_{t})\cdot\epsilon\cdot(\overline{l}_{t}^{c})^{\frac{1}{2}},$
where $\delta_{t}^{c}$ denotes the threshold of category $c$ at time step t, $\overline{l}_{t}^{c}$ is the mean predicted scores of the category $c$ at time step $t$, $\beta_{t}$ represents the update rate, and $\epsilon$ provides a linear projection. Furthermore, the threshold $\delta_{t}^{c}$ will not change if class $c$ does not exist in prediction $\hat{y}_{t}$, and we set a fixed upper and lower bound $\delta_{max}$ and $\delta_{mini}$. For unlabeled data from dynamic environments, this mechanism effectively generates an appropriate threshold for each category.
3.2.4. Data-Driven Stochastic Restoration
Existing methods address catastrophic forgetting through stochastic reset mechanisms (Wang et al., 2022) or pure data-driven restoration (Brahma and Rai, 2023). Nevertheless, the former may reset valuable parameters, potentially erasing essential knowledge relevant to the current domain. On the other hand, the latter restores the most unimportant parameters, which may retain noise parameters, thereby leading to error accumulation. Therefore, we propose the DSR strategy to reset irrelevant parameters with higher possibilities, enabling robust adaptation. Following (Brahma and Rai, 2023), we adopt the diagonal approximation of Fisher Information Matrix (FIM) (Kirkpatrick et al., 2017) based on gradient, as a metric of parameter importance. The FIM approximates parameter importance by considering the sensitivity of the loss function to parameter changes (Kirkpatrick et al., 2017). Parameters with higher values in the FIM are those for which small changes result in significant increases in loss, indicating their importance for the domain at hand.
Specifically, consider a convolution layer with weight $W_{t-1}$ and its corresponding gradient matrix after gradient descent $G_{t}$ with the same shape as $W_{t-1}$ at time step $t$. First, we use a random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$ with the same shape as the weights matrix, where each element follows a uniform distribution between 0 and 1. Moreover, we utilize the element-wise multiplication of the square of the gradient and the random matrix to approximate the FIM: $F_{t}=G_{t}\odot G_{t}\odot R_{t}$, where $\odot$ represents the element-wise multiplication. Furthermore, the proposed DSR updates the parameters based on the FIM values, which is formulated as:
(9) $M_{t}=F_{t}<\eta,$
(10) $W_{t}=M_{t}\odot W_{0}+(1-M_{t})\odot W_{t-1},$
where $M_{t}$ indicates the mask matrix of the same shape as $W_{t}$, $<$ represents the element-wise less than operation, and $\eta$ is the threshold value which is acquired by the $q$-quantile of $F_{t}$, i.e. $\eta=quantile(F_{t},q)$. Consequently, the elements in $M_{t}$ are set to 1, when the corresponding FIM value is less than $\eta$, indicating the corresponding parameters should be reset to the source weight $W_{0}$. This strategy enables the model to retain essential knowledge while introducing randomness to enhance robustness to mitigate forgetting.
4. EXPERIMENTS
Time $t\xrightarrow{\hskip 368.57964pt}$ All
Condition Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast Elastic Pixelate Jpeg Mean Gain Iter.
Source 6.8 8.1 8.0 1.5 0.2 6.8 34.6 30.7 3.0 50.2 17.6 13.5 15.1 / /
CoTTA 7.8 9.0 8.9 1.8 0.3 7.1 38.4 31.1 8.6 49.6 16.2 13.1 16.0 +0.9 6k
SVDP 7.7 10.1 9.7 2.3 0.7 13.0 42.4 45.2 15.4 47.2 21.2 14.8 19.1 +4.0 6k
IRG 8.0 11.0 9.3 3.4 1.2 13.0 37.9 41.3 15.9 38.9 16.9 13.4 17.5 +2.4 6k
MemCLR 8.5 10.4 10.6 2.7 1.1 12.2 41.4 41.6 16.4 43.1 15.4 12.7 18.0 +2.9 6k
Ours-unstop 8.6 12.1 11.7 3.6 1.5 16.7 44.7 48.1 16.7 47.4 22.5 13.9 20.6 +5.5 6k
Ours 8.7 11.7 12.4 3.6 1.5 13.5 43.2 47.3 18.4 47.0 24.8 17.2 20.8 +5.7 4.7k
Table 1. Experimental results (mAP@0.5) and adaptation iterations of Cityscapes-to-Cityscapes-C short-term CTTA task. We evaluate the performance by continually adapting the source model to the twelve corruptions. “-unstop” denotes no skipping.
Time $t\xrightarrow{\hskip 368.57964pt}$
Round 1 5 10 All
Condition Fog Motion Snow Brightness Defocus Fog Motion Snow Brightness Defocus Fog Motion Snow Brightness Defocus Mean Gain Iter.
Source 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 16.4 / /
CoTTA 38.2 10.6 0.4 33.8 9.3 40.8 10.4 0.6 36.5 9.2 40.8 10.4 0.5 36.3 9.6 19.1 +2.7 25k
SVDP 36.8 8.8 0.6 43.5 11.2 45.0 15.4 3.8 47.8 19.9 41.5 16.2 4.6 44.6 20.2 25.3 +8.9 25k
IRG 37.9 9.0 0.7 44.3 12.2 45.6 17.7 6.0 46.7 21.9 37.0 16.7 7.1 38.5 20.5 25.6 +9.2 25k
MemCLR 37.7 8.9 0.8 45.5 13.5 45.0 18.1 5.1 46.8 22.4 36.8 18.5 7.5 38.4 21.8 26.0 +9.6 25k
Ours-unstop 39.0 10.4 0.8 48.0 13.8 49.4 18.6 7.7 51.7 25.0 45.7 20.0 11.8 46.4 25.8 29.0 +12.6 25k
Ours 39.2 9.9 0.8 46.7 13.8 49.4 18.5 8.0 52.9 24.7 46.2 19.5 11.3 47.3 29.1 29.2 +12.8 20.1k
Table 2. Experimental results (mAP@0.5) and adaptation iterations of Cityscapes-to-Cityscapes-C long-term CTTA task. We evaluate the performance by continually adapting the source model to the five corruption ten times. To save space, we display selected rounds of results, and full results are provided in the appendix. “-unstop” denotes no skipping.
In this study, we rigorously evaluate our methodology across four benchmark tasks in CTTA object detection. These tasks encompass continual adaptation to synthetic and real-world distribution shifts, evaluated over short and long-term periods. Inspired by the foundational work CoTTA (Wang et al., 2022), the short-term CTTA task entails the sequential adaptation to various target domains once. In contrast, the long-term CTTA task involves continually adapting the model toward groups of target domains cyclically.
4.1. Datasets
Cityscapes. The Cityscapes dataset (Cordts et al., 2016) is collected for urban scene understanding, encompassing 2,975 training images and 500 validation images with eight object types. We utilize the model pre-trained on the training set of Cityscapes as the source model, and the source data is discarded during adaptation.
Cityscapes-C. The Cityscapes-C (Hendrycks and Dietterich, 2019) is initially designed to assess robustness against various corruptions, introducing 15 types of corruption with 5 levels of severity. We create the dataset by applying these corruptions at the maximum severity level 5 to the validation set of clean cityscapes, treating each corruption type as an individual target domain comprising 500 images.
Our short-term CTTA task selectively focuses on the latter 12 corruptions. For the long-term CTTA task, we prioritize the five corruptions related to autonomous driving scenarios as the target domain groups following (Gan et al., 2023), namely Fog, Motion, Snow, Brightness, and Defocus. We repeat the adaptation to the target domain group 10 times to evaluate long-term performance.
SHIFT. The SHIFT (Sun et al., 2022) is a synthetic dataset for autonomous driving, featuring real-world environmental changes. SHIFT can be categorized as clear, cloudy, overcast, rainy, and foggy, where each condition contains images taken at various times ranging from daytime to night. For SHIFT, short-term CTTA tasks are considered. We designate the clear condition as the source domain, with the remaining four conditions as the target domain groups including nearly 20k images in total.
ACDC. The ACDC dataset (Sakaridis et al., 2021) shares the same class types as Cityscapes and is collected in four different adverse visual conditions, including Fog, Night, Rain, and Snow. Following (Wang et al., 2022), we use these four conditions as the target domain group for the long-term CTTA task, with 400 unlabeled images per condition. Similarly, the source-trained model is continually adapted to the target domain group for 10 cycles.
4.2. Implementation Details
We adopt the Faster R-CNN with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) as the backbone. Aligning with (VS et al., 2023b; Wang et al., 2022), we maintain a batch size of 1, thereby emulating a real-world application scenario where the detection model adapts toward a continuous influx of images. The source models are trained using an SGD optimizer with a learning rate of 0.001 and a momentum of 0.9. Algorithms are implemented leveraging the Detectron2 platform (Wu et al., 2019). The metric of mAP at an IoU threshold of 0.5 is employed for evaluation. Each experiment is conducted on 1 NVIDIA A800 GPU. More hyper-parameters and implementation details are provided in the appendix.
4.3. Baselines and Compared Approaches
We compare CTAOD with six baselines across various settings, including Source (Ren et al., 2015), IRG (VS et al., 2023a), Memclr (VS et al., 2023b), CoTTA (Wang et al., 2022), SVDP (Yang et al., 2024), and CTAOD-upstop. Specifically, ”Source” represents the source model without adaptation. Memclr and IRG are SFDA object detection methods, also referred to as TTA methods (Wang et al., 2022). MemCLR integrates cross-attention with CL, while IRG incorporates instance relation graph and CL. CoTTA utilizes weight-averaged predictions and random neurons recovery to tackle CTTA, while SVDP explores sparse visual prompts for CTTA dense prediction. “CTAOD-unstop” represents CTAOD without dynamic skipping.
4.4. Results and Analysis
Time $t\xrightarrow{\hskip 368.57964pt}$
Round 1 4 7 10 All
Condition Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Mean Gain Iter.
Source 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 36.0 / /
CoTTA 53.7 19.7 38.0 42.4 53.1 19.7 37.7 42.9 51.3 19.0 36.5 41.8 50.9 19.1 36.0 42.4 37.8 +1.8 16k
SVDP 52.8 20.0 35.6 42.0 54.6 23.5 38.7 43.8 52.8 24.0 38.6 43.9 51.8 23.6 38.2 43.0 39.5 +3.5 16k
IRG 52.7 20.6 36.0 42.9 53.6 23.2 38.1 45.6 51.0 23.1 37.1 42.8 49.7 22.5 36.5 40.7 38.9 +2.9 16k
MemCLR 52.9 21.2 35.2 42.8 52.8 23.2 38.2 44.4 51.4 23.2 37.1 42.9 49.9 22.9 36.6 41.0 38.7 +2.7 16k
Ours-unstop 52.9 20.4 34.5 42.9 54.5 23.8 39.6 45.8 53.5 25.1 39.0 45.3 52.5 24.3 38.8 44.7 40.3 +4.3 16k
Ours 52.8 20.1 35.2 42.9 54.4 24.6 38.6 45.5 53.5 25.6 38.5 46.2 53.5 25.4 37.4 45.3 40.4 +4.4 13.2k
Table 3. Experimental results (mAP@0.5) and adaptation iterations of Cityscapes-to-ACDC long-term CTTA task. We evaluate the performance by continually adapting the source model to the four conditions ten times. To save space, we display selected rounds of results, and full results are provided in the appendix. “-unstop” denotes no skipping.
(a) Source
(b) MemCLR
(c) CTAOD (ours)
Figure 3. Qualitative results. We compare the detection results of the Source model, MemCLR, and CTAOD in the $10^{th}$ round of adaption to defocus corruption on the long-term Cityscapes-to-Cityscapes-C task.
4.4.1. Synthetic Continual Distribution Shift
We first evaluate the effectiveness of CTAOD on the short-term Cityscapes-to-Cityscapes-C adaptation tasks. As depicted in Table 1, CoTTA, IRG, and Memclr exhibit enhancements in performance, achieving 16.0, 17.5, and 18.0 mAP respectively. Although the SFDA methods like IRG and MemCLR assume a static target domain, employing a large momentum update rate $\alpha$ for the teacher enables a relatively stable adaptation in the short term. Consequently, CL leads to a better performance than CoTTA which solely employs weight-averaged predictions. Moreover, SVDP, which utilizes visual prompts, achieves a sub-optimal performance of 19.1 mAP. Remarkably, our proposed method improves the performance to 20.8 mAP, consistently outperforming the above approaches. Additionally, the dynamic skipping strategy in AM reduces adaptation iterations by 20% while achieving a 0.2 mAP improvement compared to CTAOD-unstop. CTAOD ensures a more reliable and efficient adaptation to target domains characterized by intense and continual changes in the short term.
As presented in Table 8, the long-term task outcomes reveal the source model’s poor performance, with an average mAP of 16.4. Despite the improvement of Memclr and IRG, their performance also begins to decline in the later rounds, failing to maintain stability over long-term adaptation. We believe this is due to the aforementioned methods not accounting for continual distribution shifts, resulting in error accumulation and catastrophic forgetting. Furthermore, CoTTA utilizes a stochastic restoration mechanism to mitigate forgetting, but its randomness may result in losing crucial information, thus limiting its performance at 19.1 mAP. In contrast, SVDP which employs sparse visual prompts raises the performance to 25.3 mAP. Particularly, CTAOD yields a remarkable 12.8 mAP enhancement over the Source and surpasses all comparative baselines using 80% adaptation iterations, which employs the DSR mechanism to conserve valuable knowledge while eliminating noise from prior domains. These findings empirically validate the effectiveness of our proposed method in ensuring stable and efficient adaptation amidst synthetic continual distribution shifts, across both short-term and long-term scenarios.
Time $t\xrightarrow{\hskip 195.12767pt}$ All
Condition Cloudy Overcast Rainy Foggy Mean Gain Iter.
Source 51.8 41.5 43.8 33.9 42.7 / /
CoTTA 51.1 40.1 40.7 29.9 40.5 -2.2 19.6k
SVDP 52.0 41.0 43.8 35.9 43.2 +0.5 19.6k
IRG 51.9 40.6 42.7 34.3 42.4 -0.3 19.6k
MemCLR 51.8 40.4 42.6 35 42.4 -0.3 19.6k
Ours-unstop 52.3 41.3 44.1 36.8 43.6 +0.9 19.6k
Ours 52.3 41.6 44.4 37.3 43.9 +1.2 16.4k
Table 4. Experimental results (mAP@0.5) of SHIFT short-term CTTA task. We evaluate the performance by continually adapting the source model to the four conditions.
4.4.2. Real-World Continual Distribution Shift
We also evaluate our method on the real-world continual distribution shift dataset. The results of the SHIFT short-term CTTA task are shown in Table 4. While other baseline methods suffer from performance decline, SVDP which employs a fixed threshold for pseudo-labeling and prompts learning achieves a 0.5 mAP improvement. Moreover, CTAOD which introduces a dynamic threshold strategy in AM shows a superior performance of 43.9 mAP with 16% improvement in efficiency.
The experimental results of ACDC long-term CTTA task are summarized in Table 3. CoTTA, IRG, and MemCLR potentially suffer from error accumulation and catastrophic forgetting, manifesting in a rapid decline in performance during the later stages. Similarly, SVDP achieves a sub-optimal performance of 39.5 mAP. Notably, CTAOD achieves a performance of 40.4 mAP, yielding an absolute improvement of 0.9 mAP over baselines. While CTAOD may not display a significant advantage over other methods in the initial round, its performance continually improves and remains stable throughout the long-term adaptation. These results underscore the capability of our method to foster robust adaptation toward the real-world continual distribution shift in both the short and long term.
4.4.3. Qualitative Results
As shown in Figure 3, we selectively compare the source model, MemCLR, and CTAOD in the final round adaptation to the brightness on the long-term Cityscapes-to-Cityscapes-C tasks. CTAOD guides the model in learning better feature representations while effectively mitigating forgetting in the long-term adaptation. Therefore, CTAOD assists the detector in distinguishing more foreground object categories and better locating them. More visualization analysis is presented in the appendix.
Mean-Teacher OCL AM DSR Mean Gain
1 $\checkmark$ 25.9 /
2 $\checkmark$ $\checkmark$ 26.7 +0.8
3 $\checkmark$ $\checkmark$ 27.3 +1.4
4 $\checkmark$ $\checkmark$ 27.4 +1.5
5 $\checkmark$ $\checkmark$ $\checkmark$ 28.6 +2.7
6 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 29.2 +3.3
Table 5. Alation experiment. “Mean-Teacher” represents the base mean teacher framework with weak-strong augmentation and KL divergence distillation. All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
4.4.4. Ablation Study
We conduct an ablation study to empirically assess the impact of the OCL, AM, and DSR. As shown in Table 5, the integration of the mean-teacher architecture elevates the performance to 25.9 mAP. Building upon the mean-teacher foundation, the proposed OCL, AM, and DSR achieve an additional 1.5, 1.4, and 0.8 mAP improvement individually. Moreover, the combination of the OCL and AM modules raises the performance to 28.6 mAP. Furthermore, the inclusion of the DSR mechanism brings a 0.6 mAP boost in performance, culminating in a comprehensive model performance of 29.2 mAP. More fine-grained comparisons with AM and DSR are provided in the appendix.
5. CONCLUSION
In this work, we proposed CTAOD, aimed at addressing the challenges of low-quality pseudo-labels and random errors resulting from stochastic recovery in continual test-tiem adaptation. Firstly, object-level contrastive learning leverages teacher ROI features for contrastive learning to refine the feature representation, tailored for object detection. Secondly, the dynamic threshold strategy selects higher-quality pseudo-labels by dynamically updating the category-specific threshold based on the predicted confidence scores. Lastly, the data-driven stochastic restoration mechanism selectively reset inactive parameters to mitigate forgetting while retaining valuable knowledge. The empirical results of four continual test-time adaptation tasks for object detection demonstrate the efficacy of CTAOD in both short-term and long-term adaption scenarios.
Appendix A Appendix Overview
In this appendix, we provide additional details and analysis. In the first section, we provide implementation details on CTAOD and other baseline methodologies. In the second section C, we furnish complete results for the quantitative research and additional visualization examples for the qualitative research. In the third section, we further discuss the continual learning. In the last section E, we discuss the limitations and prospective directions of our research.
Appendix B Additional Implementation Details
Following (VS et al., 2023b), we trained a Faster R-CNN (Ren et al., 2015) with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) on the training set of Cityscapes (Cordts et al., 2016) as the source-trained model of Cityscapes. We utilized an SGD optimizer with a learning rate of 0.001, a momentum of 0.9, and a batch size of 1 and 70k iterations. Similarly, we trained another source model on the clear condition of the SHIFT (Sun et al., 2022) with 30 epochs. We will specify the implementation details of the CTAOD and other baseline methods in the following subsections.
B.1. Implememtation of CTAOD
Detailed hyper-parameters of CTAOD for each benchmark are listed in Table 7. Furthermore, we present a pseudo-code outline of our CTAOD framework in Algorithm 1. The three key components of CTAOD are highlighted, i.e., Object-level Contrastive Learning, Adaptive Monitoring, and Data-driven Stochastic Restoration.
B.2. Implememtation of Baseline Methods
The CoTTA (Wang et al., 2022) and SVDP (Yang et al., 2024) are originally designed for the classification or segmentation tasks. We implement them based on their public code, with some modifications to fit into the objection detection. Specifically, we implement the Cotta method by constructing a mean-teacher framework based on Faster R-CNN, which incorporates a stochastic restoration mechanism. Similarly, we implement SVDP with visual prompts on the area of the bounding box in pseudo-labels based on mean-teacher framework. Further, the MemCLR (VS et al., 2023b) and IRG (VS et al., 2023a) are source-free domain adaptation methods for object detection. Therefore, we implement the MemCLR and IRG based on their public code, with little modifications to follow the CTTA setting (i.e., updating the model parameters during evaluation). For a fair comparison, we set the same momentum $\alpha$ of the Exponential Moving Average (EMA) for CoTTA, SVDP, MemCLR, and IRG as CTAOD, all of which apply the mean-teacher framework (Tarvainen and Valpola, 2017). Other hyper-parameters are retained as specified in their public code.
Appendix C Additional Experiment Results
C.1. Quantitative Results
We present the complete experiment results on the two long-term CTTA tasks, i.e., Cityscapes-to-Cityscapes-C and Cityscapes-to-ACDC adaptation tasks, in Table 8 and 9. Experiments show that the proposed CTAOD can largely maintain strong performance in the long term.
C.2. Qualitative Results
As shown in figure 4 - 8, we present more visualization results to compare the CTAOD and other baseline methods based on the final round adaption on the long-term Cityscapes-to-Cityscapes-C task. The results further validate the effectiveness of CTAOD in distinguishing more foreground object categories and better locating them, compared with other baseline methods.
C.3. Additional Ablation Study
CTAOD-unstop FT of 0.9 FT of 0.8 FT of 0.7 SR DR Mean
1 $\checkmark$ $\checkmark$ 27.7
2 $\checkmark$ $\checkmark$ 28.2
3 $\checkmark$ $\checkmark$ 28.5
4 $\checkmark$ $\checkmark$ 27.1
5 $\checkmark$ $\checkmark$ 28.3
6 $\checkmark$ 29.0
Table 6. Additional Ablation Study. The “CTAOD-unstop” denotes CTAOD without dynamic skipping. The “FT” represents replacing dynamic threshold in CTAOD with a fixed threshold. The “SR” denotes the stochastic restoration (Wang et al., 2022), and the “DR” indicates the data-driven restoration (Brahma and Rai, 2023). All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
As presented in Table 6, we perform an additional ablation study to evaluate the effectiveness of the proposed Dynamic Threshold (DT) strategy in Adaptive Monitoring and Data-driven Stochastic Restoration (DSR). The findings indicate that the DT approach results in improvements of 1.3, 0.8, and 0.5 mAP over fixed thresholds set at 0.9, 0.8, and 0.7, respectively. While the stochastic restoration and data-driven restoration limit the performance to 27.1 and 28.3 mAP, the DSR mechanism raises the performance to 29.0 mAP.
Hyper-parameter Description City2City-C-Short City2City-C-Long SHIFT-Short City2ACDC-Long
$\gamma$ Learning rate 0.001 0.001 0.0001 0.001
$\alpha$ Momentum in EMA 0.9996 0.9998 0.9999 0.9998
$\tau$ Temperature in CL 0.07 0.07 0.07 0.07
$\delta_{0}$ Initial valuse in DT 0.8 0.8 0.8 0.8
$\beta_{t}$ Hyper-parameter in DT 0.95 0.95 0.95 0.95
$\epsilon$ Hyper-parameter in DT 1.3 1.3 1.3 1.3
$\delta_{max}$ The upper-bound in DT 0.9 0.9 0.9 0.9
$\delta_{mini}$ The lower-bound in DT 0.7 0.7 0.7 0.7
$\beta_{s}$ Hyper-parameter in DS 0.7 0.75 0.6 0.7
$\delta_{s}$ The threshold in DS 1.5 1.4 1.4 1.4
$q$ Quantile value 0.01 0.001 0.001 0.01
Table 7. Detailed hyper-parameters for each benchmark. “DT” and “DS” denotes dynamic threshold and dynamic skipping in adaptive monitoring. “City2City-C-Short” denotes the short-term Cityscapes-to-Cityscapes-C task, “City2City-C-Long” denotes long-term Cityscapes-to-Cityscapes-C task, and “ City2ACDC-Long” denotes long-term Cityscapes-to-ACDC task.
Time $t\xrightarrow{\hskip 403.26341pt}$
Round 1 2 3 4 5 /
Condition F. M. S. B. D. F. M. S. B. D. F. M. S. B. D. F. M. S. B. D. F. M. S. B. D. / /
Source 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 / /
CoTTA 38.2 10.6 0.4 33.8 9.3 39.9 10.7 0.5 34.0 9.9 40.6 10.4 0.6 34.4 9.8 40.9 10.4 0.6 35.4 9.1 40.8 10.4 0.6 36.5 9.2 / /
SVDP 36.8 8.8 0.6 43.5 11.2 43.9 11.8 1.7 47.9 14.9 46.2 13.9 2.5 50.0 17.8 46.2 14.4 3.2 49.0 19.6 45.0 15.4 3.8 47.8 19.9 / /
IRG 37.9 9.0 0.7 44.3 12.2 46.1 12.9 2.0 49.2 16.6 47.6 15.2 3.0 49.5 19.5 45.2 17.3 4.6 48.6 21.4 45.6 17.7 6.0 46.7 21.9 / /
MemCLR 37.7 8.9 0.8 45.5 13.5 45.8 13.6 2.3 49.8 17.7 47.3 16.4 4.1 50.2 20.3 46.4 17.4 4.8 47.8 21.9 45.0 18.1 5.1 46.8 22.4 / /
Ours-unstop 39.0 10.4 0.8 48.0 13.8 48.2 13.7 3.1 52.9 18.1 49.3 15.9 5.3 53.3 21.3 48.5 17.1 7.1 52.2 23.5 49.4 18.6 7.7 51.7 25.0 / /
Ours 39.2 9.9 0.8 46.7 12.7 46.8 13.5 2.7 52.3 17.9 49.1 15.8 4.5 53.9 21.2 48.8 17.7 6.9 53.8 23.1 49.4 18.5 8.0 52.9 24.7 / /
Round 6 7 8 9 10 Mean Gain
Source 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 16.4 /
CoTTA 39.6 10.3 0.6 34.9 9.4 40.2 10.8 0.5 35.3 10.0 40.0 10.8 0.5 35 9.7 40.1 10.5 0.5 34.1 9.5 40.8 10.4 0.5 36.3 9.6 19.1 +2.7
SVDP 44.4 15.6 4.2 47.6 20.5 44.1 16.1 4.3 46.8 20.4 43.0 15.9 4.6 45.7 20.7 42.6 16.0 4.8 45.0 20.2 41.5 16.2 4.6 44.6 20.2 25.3 +8.9
IRG 43.9 18 5.9 44.8 22.1 42.3 17.5 6.6 42.9 22.2 41.7 16.9 7.2 42.1 21.8 38.8 17 7.2 40.1 20.6 37.0 16.7 7.1 38.5 20.5 25.6 +9.2
MemCLR 43.4 19.3 5.9 44.6 23.3 41.1 19.5 6.6 43.1 23.9 39.6 19.1 7.0 40.9 24.3 38.6 18.9 7.2 39.1 22.4 36.8 18.5 7.5 38.4 21.8 26.0 +9.6
Ours-unstop 48.5 19.7 8.9 50.7 26.3 47.8 19.6 9.8 49.0 26.1 45.7 19.3 10.5 49.4 26.9 45.2 19.7 11.2 47.4 26.9 45.7 20.0 11.8 46.4 25.8 29.0 +12.6
Ours 48.6 19.8 9.1 51.2 25.4 48.8 20.1 10.0 49.5 27.0 47.3 19.9 10.0 49.2 27.9 46.9 19.4 11.0 47.5 28.9 46.2 19.5 11.3 47.3 29.1 29.2 +12.8
Table 8. Complete experimental results (mAP@0.5) of Cityscapes-to-Cityscapes-C long-term CTTA task. We evaluate the long-term adaptation performance by continually adapting the source model to the five corruption ten times with the largest corruption severity level 5. “F.”, “M.”, “S.”, “B.”, and “D.” represent Fog, Motion, Snow, Brightness, and Defocus, respectively.
Appendix D Additional Related Work
D.1. Continual Learning
Continual learning, also known as incremental learning or life-long learning, typically involves enabling the model to retain previously acquired knowledge while learning from a sequential series of tasks, i.e. preventing catastrophic forgetting (De Lange et al., 2021; Parisi et al., 2019). It is commonly categorized into replay methods (Rebuffi et al., 2017; Tiwari et al., 2022; Rolnick et al., 2019), parameter isolation method (Aljundi et al., 2017; Xu and Zhu, 2018), and regularization-based methods (Kirkpatrick et al., 2017; Li and Hoiem, 2017; Zenke et al., 2017). As an illustration, Elastic weight consolidation (EWC) (Kirkpatrick et al., 2017) is a regularization-based technique that penalizes the changing of parameters with a significant impact on prediction, based on the Fisher Information Matrix (FIM). In this paper, motivated by (Kirkpatrick et al., 2017; Brahma and Rai, 2023), we utilize gradients to approximate FIM as a metric of parameter importance for resetting noisy parameters. Furthermore, we introduce randomness to enhance model robustness during continual adaptation. Additionally, while the continual learning approaches aim to tackle catastrophic forgetting in sequences of new tasks, our work focuses on learning from different domains for a single task.
Appendix E Limitation
Firstly, our experiments are primarily based on a single detector backbone, i.e. Faster R-CNN (Ren et al., 2015). Future work will extend to other advanced detector backbone networks to assess the universality and generalization capabilities of CTAOD. Moreover, the proposed DSR mechanism to reset the model every iteration increases the computational cost. Future work could explore more efficient mechanisms for restoration, such as restoring the model only when significant changes in the distribution of the target domain are detected. Finally, the evaluation tasks in our work are designed to simulate real-world adaptation scenarios by incorporating datasets affected by corruption and adverse conditions. However, real-world data distributions are inherently more complex. Consequently, a promising avenue for future research would be to apply our methodology in practical, real-world systems to further validate its effectiveness.
Time $t\xrightarrow{\hskip 390.25534pt}$
Round 1 2 3 4 5 /
Condition Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow / /
Source 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 / /
CoTTA 53.7 19.7 38 42.4 53.6 20.3 37.7 42.2 53.7 20.1 38.1 43.3 53.1 19.7 37.7 42.9 52.6 19.5 37 42.2 / /
SVDP 52.8 20.0 35.6 42.0 53.4 22.0 37.4 43.8 54.0 23.0 38.8 44.0 54.6 23.5 38.7 43.8 53.4 24.0 38.0 44.4 / /
IRG 52.7 20.6 36 42.9 53.2 22.4 37.4 44.6 53.8 23.2 38.1 45.5 53.6 23.2 38.1 45.6 53.5 23.2 38.1 45.3 / /
MemCLR 52.9 21.2 35.2 42.8 52.5 23.1 37.3 44.3 53.9 23.2 38.4 44.8 52.8 23.2 38.2 44.4 51.5 23.5 37.2 43.8 / /
Ours-unstop 52.9 20.4 34.5 42.9 53 22.9 38.6 44.6 54.2 24 39.5 45.3 54.5 23.8 39.6 45.8 54.4 24.4 39.2 45.4 / /
Ours 52.8 20.1 35.2 42.9 53.6 22.2 38.3 44.2 53.9 23.5 39.6 45.0 54.4 24.6 38.6 45.5 54.6 25.1 39.2 46.3 / /
Round 6 7 8 9 10 Mean Gain
Source 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 36 /
CoTTA 52.3 19.1 36.3 42 51.3 19 36.5 41.8 51.2 18.9 36.1 41.4 51.5 19.2 36.4 41.3 50.9 19.1 36 42.4 37.8 +1.8
SVDP 53.3 24.0 38.9 43.8 52.8 24.0 38.6 43.9 52.0 23.8 38.7 43.4 51.8 23.8 38.6 43.6 51.8 23.6 38.2 43.0 39.5 +3.6
IRG 52.6 23.5 37.5 43.7 51 23.1 37.1 42.8 50.2 23.1 37 42.2 49.8 23.2 36.5 41.1 49.7 22.5 36.5 40.7 38.9 +2.9
MemCLR 50.2 23.3 36.6 43 51.4 23.2 37.1 42.9 50.5 23.1 36.6 42.6 49.9 22.8 36.8 42 49.9 22.9 36.6 41 38.7 +2.7
Ours-unstop 53.9 24.9 39.2 45.1 53.5 25.1 39 45.3 53.3 24.6 39.3 45.3 53.2 24.5 39.3 44.9 52.5 24.3 38.8 44.7 40.3 +4.3
Ours 54.4 25.4 38.7 46.1 53.5 25.6 38.5 46.2 54.1 25.9 38.0 45.8 53.3 25.9 37.9 45.4 53.5 25.4 37.4 45.3 40.4 +4.4
Table 9. Complete experimental results (mAP@0.5) of Cityscapes-to-ACDC long-term CTTA task. We evaluate the long-term adaptation performance by continually adapting the source model to the four conditions ten times with the largest corruption severity level 5.
Algorithm 1 Pseudo-code for the CTAOD Method
1:Initialization: Object detectors: The student and teacher model parameters $\theta_{0}^{std}$ and $\theta_{0}^{tch}$ initialized from source-trained model $\theta_{0}$. Hyper-parameters: learning rates $\gamma$, Momentum $\alpha$ in exponential moving average (EMA), temperature $\tau$, thresholds $\delta_{0}^{c}$, $\beta_{t}$ and $\epsilon$ in dynamic threshold, $\beta_{s}$ and $\delta_{s}$ in dynamic skipping, and quantile value $q$.
2:Input: Unlabeled test data $x_{t}$ for the target domain
3:for each iteration $t$ do
4:       Generate weakly and strongly augmented images $f_{weak}(x_{t})$ and $f_{strong}(x_{t})$
5:       Generate predictions: $y_{t}=f_{\theta_{t-1}^{tch}}(f_{weak}(x_{t}))$ and proposal $p_{t}$ using the teacher model
6:       // 1. Adaptive Monitoring
7:       Compute the mean predicted scores of all instances $\overline{l}_{t}$ and the index $\frac{\overline{l}_{t}}{\overline{l}_{t}^{ema}}$
8:       Udapate the exponentially moving average of mean scores $\overline{l}_{t}^{ema}$ through:
9:             $\overline{l}_{t}^{ema}\leftarrow\beta_{s}\cdot\overline{l}_{t-1}^{ema}+(1-\beta_{s})\cdot\overline{l}_{t},$
10:       if $\frac{\overline{l}_{t}}{\overline{l}_{t}^{ema}}$ not lies with $(\frac{1}{\delta_{s}},\delta_{s})$ then
11:             Pause the current adaption
12:       end if
13:       for each category $c$ do
14:             Compute the mean predicted scores $\overline{l_{t}^{c}}$ of category $c$
15:             Update dynamic thresholds based on predictions:
16:                   $\delta_{t}^{c}\leftarrow\beta_{t}\cdot\delta_{t-1}^{c}+(1-\beta_{t})\cdot\epsilon\cdot(\overline{l_{t}^{c}})^{\frac{1}{2}}$
17:       end for
18:       // 2. Object-level Contrastive Learning
19:       Extract teacher features $F_{t}^{T}$ and student features $F_{t}^{S}$ based on $p_{t}$ using RoIAlign
20:       Compute contrastive learning loss $\mathcal{L}_{cl}$:
21:             $\mathcal{L}_{cl}(x_{t})=\frac{1}{l}\sum_{i=1}^{l}-\log\frac{\exp(f_{i}^{T}\cdot f_{i}^{S}/\tau)}{\sum_{j=1}^{l}\exp(f_{i}^{T}\cdot f_{j}^{S}/\tau)}$
22:       Generate the pseudo-label $\hat{y}_{t}$ through the category-specific threshold $\delta_{0}^{c}$
23:       Compute the supervised loss $\mathcal{L}_{pl}(x_{t})=\mathcal{L}_{rpn}(x_{t},\hat{y}_{t})+\mathcal{L}_{rcnn}(x_{t},\hat{y}_{t})$
24:       Perform gradient descent to update the student model $\theta_{t}^{std}$ and get the gradient matrix $G_{t}$
25:       Update teacher model $\theta_{t}^{tch}$ via EMA strategy
26:       // 3. Data-driven Stochastic Restoration
27:       Generate random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$
28:       Generate the fisher information matrix: $F_{t}=G_{t}\odot G_{t}\odot R_{t}$
29:       Find the $q$-quantile of $F_{t}$: $\eta=quantile(F_{t},q)$
30:       Generate the mask matrix: $M_{t}=F_{t}<\eta$
31:       Reset the updated student model: $\theta_{t}^{std}=M_{t}\odot\theta_{0}^{std}+(1-M_{t})\odot\theta_{t}^{std}$
32:end for
33:Output: Teacher predictions $y_{t}$
(a) Source
(b) SVDP
(c) CoTTA
(d) IRG
(e) MemCLR
(f) CTAOD (ours)
Figure 4. Complete qualitative results. We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Brightness corruption on the long-term Cityscapes-to-Cityscapes-C task.
(a) Source
(b) SVDP
(c) CoTTA
(d) IRG
(e) MemCLR
(f) CTAOD (ours)
Figure 5. Complete qualitative results. We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Defocus corruption on the long-term Cityscapes-to-Cityscapes-C task.
(a) Source
(b) SVDP
(c) CoTTA
(d) IRG
(e) MemCLR
(f) CTAOD (ours)
Figure 6. Complete qualitative results. We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Fog corruption on the long-term Cityscapes-to-Cityscapes-C task.
(a) Source
(b) SVDP
(c) CoTTA
(d) IRG
(e) MemCLR
(f) CTAOD (ours)
Figure 7. Complete qualitative results. We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Motion corruption on the long-term Cityscapes-to-Cityscapes-C task.
(a) Source
(b) SVDP
(c) CoTTA
(d) IRG
(e) MemCLR
(f) CTAOD (ours)
Figure 8. Complete qualitative results. We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Snow corruption on the long-term Cityscapes-to-Cityscapes-C task.